OpenStudy (anonymous):

Solve using the Quadratic Formula. 2x² - 6 = 0

OpenStudy (anonymous):

Do you really need to? You can simply add 6 to both sides divide by 2 and then square root both sides.

OpenStudy (anonymous):

yes I really need to

OpenStudy (anonymous):

that sucks well can you tell me what a b and c are?

OpenStudy (zepp):

\[\large \begin{align} 2x^2 - 6 &= 0\\ 2x^2&=6\\ x^2&=6\div2 \\ x^2&=3\\ x&=\pm\sqrt{3} \end{align}\]

OpenStudy (zepp):

I see no need to use the quadratic formula here :$

OpenStudy (anonymous):

me too but if the questions asks you to do it then you have no choice

OpenStudy (anonymous):

why do I not need to use it.

OpenStudy (zepp):

because you can solve it algebraically

OpenStudy (anonymous):

can you tell me what a b and c are? I know you have done problems like this.

OpenStudy (anonymous):

I'm so bad at math but here I go x^2-2x-6 so b=2,c=6

OpenStudy (zepp):

A wild -2x appeared, the leading coefficient disappeared..

OpenStudy (zepp):

\(\large x^2-2x-6=0\) A should be 1 B should be -2 C should be -6

OpenStudy (anonymous):

ok

OpenStudy (zepp):

A quadratic equation has the form \(\large ax^2+bx+c=0\), where a, b and c are some constants. Your question is \(\large x^2−2x−6=0\), we can rewrite that as \(\large (1)x^2+(-2)x+(-6)=0\)

OpenStudy (zepp):

That's why a=1; b=-2 and c=-6

OpenStudy (zepp):

Notice the +/- sign, they are very important Now use the quadratic formula, which is \(\huge x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\), simply plug in the values (BE CAREFUL WITH SIGNS!! I would suggest you to put parenthesis) \(\huge x=\frac{-(-2)\pm\sqrt{(-2)^2-4(1)(-6)}}{2(1)}\) And the rest is arithmetic :P

OpenStudy (zepp):

So the first difficult part is to identify what's a,b and c (just be careful with the signs and you'll be fine) then simply plug in the values in and you're done!

OpenStudy (anonymous):

Thank you @zepp

OpenStudy (zepp):

\(\huge x=\frac{-(-2)\pm\sqrt{(-2)^2-4(1)(-6)}}{2(1)}\\\huge \frac{2\pm\sqrt{4+24}}{2}=\frac{2\pm\sqrt{28}}{2}\) There you :D

OpenStudy (zepp):

You are welcome :)

hero (hero):

If it says "use the quadratic formula" you have to use it. If you use an alternative method, you will get partial credit. If you receive a warning from the teacher about it and continue to do it, you will lose more points. If you use a method other than the one asked of you on a test, you will receive zero points.

OpenStudy (zepp):

^